A consumer response team hears directly from consumers about the challenges they face in the​ marketplace, brings their concerns to the attention of financial​ institutions, and assists in addressing their complaints. The consumer response team accepts complaints related to​ mortgages, bank accounts and​ services, private student​ loans, other consumer​ loans, and credit reporting. An analysis of complaints over time indicates that the mean number of credit reporting complaints registered by consumers is 2.35 per day. Assume that the number of credit reporting complaints registered by consumers is distributed as a Poisson random variable

what is the probability that one credit complaint will be regisered by consumers
Answer #1

Let X be the random variable denoting the number of credit reporting complaints registered by the consumers.

Thus, X ~ Poi(2.35)

Probability ,P(X=x) = P(x; μ) = (e) (μx) / x!

the probability that one credit complaint will be registered by consumers

μ = 2.35

P(X=1) = (e-2.35) (2.351) / 1! = 0.2241

*Please ask if you have any doubts,happy to help you.

*Please give me up vote, if it is useful to you,Thank you.